OpenStudy (anonymous):

How can I find critical number of f(x)=x/(x^2-x+1) [0,3]. When I solve 3x^2-2x+1 , it is complex number, so, I am confused about the result.

OpenStudy (anonymous):

to find the critical cnumber, you first obtain its derivative

OpenStudy (anonymous):

yes, when I solve the derivative , it is not real numbers

OpenStudy (anonymous):

\[ f(x)=\frac{x}{x^2-x+1}\\ f'(x)=\frac{1(x^2-x+1)-(x)(2x-1)}{(x^2-x+1)}=0\\ \implies x^2-x+1-2x^2+x=0\\ \implies -x^2+1=0\implies x^2-1=0\\ \implies (x-1)(x+1)=0 \]

OpenStudy (anonymous):

@Yating can you take it from there?

OpenStudy (anonymous):

Yes, thanks, I did differently from what you showed. Let me check where I made mistakes

OpenStudy (anonymous):

Oh, I misused quotient rule. Thanks a lot.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

also, I missed to put the square part in the denominator of the quotient rule

OpenStudy (anonymous):

I figured out, but the most important part in this case is numerator.